x/3 - 1/3 = 2/x
What is x?

Answer by mukhopadhyay(490)

x/3 - 1/3 = 2/x
=> (x-1)/3 = 2/x
=> x(x-1) = 6
=> x^2-x-6 = 0
=> (x-3)(x+2) = 0
=> x=3 or x=-2